OpenStudy (tatianagomezb):

How do I find the zeros of this function without graphing. f(x)=3+3 cos (x-π) I know the graph has lot of zeros. How would I find at least two and explain they repeat periodically? Thanks

OpenStudy (anonymous):

\[f \left( x \right)=3+3\cos (x-\pi)=3+3\cos \left\{ -\left( \pi-x \right) \right\}=3+\cos \left( \pi-x \right)=3-3\cos x\] \[3-3\cos x=0,\cos x=1=\cos \left( 2n+1 \right)\frac{ \pi }{ 2 },x=\frac{ n \pi }{ 2 }\]

OpenStudy (tatianagomezb):

3+3cos{−(π−x)}=3+cos(π−x)=3−3cosx Good start, thanks but I don't understand why is this step valid. Why does the minus sign hace to change.

OpenStudy (anonymous):

\[\cos \left( -x \right)=\cos x,\cos \left( \pi-x \right)=-\cos x\]

OpenStudy (tatianagomezb):

Okay, I got it. Thanks again

OpenStudy (anonymous):

yw